# Conceptualization of this study, Methodology, Software

### v1

@@ Abstract
Electroencephalogram (EEG) signals play a pivotal role in clinical medicine, brain research, and neurological disease studies. However, susceptibility to various physiological and environmental artifacts introduces noise in recorded EEG data, impeding accurate analysis of underlying brain activity. Denoising techniques are crucial to mitigate this challenge. Recent advancements in deep learning-based approaches exhibit substantial potential for enhancing the signal-to-noise ratio of EEG data compared to traditional methods. In the realm of large-scale language models (LLMs), the Retentive Network (Retnet) infrastructure, prevalent for some models, demonstrates robust feature extraction and global modeling capabilities. Recognizing the temporal similarities between EEG signals and natural language, we introduce the Retnet from natural language processing to EEG denoising. This integration presents a novel approach to EEG denoising, opening avenues for a profound understanding of brain activities and accurate diagnosis of neurological diseases. Nonetheless, direct application of Retnet to EEG denoising is unfeasible due to the one-dimensional nature of EEG signals, while natural language processing deals with two-dimensional data. To facilitate Retnet application to EEG denoising, we propose the signal embedding method, transforming one-dimensional EEG signals into two dimensions for use as network inputs. Experimental results validate the substantial improvement in denoising effectiveness achieved by the proposed method.
@@ 1 Introduction
Electroencephalography (EEG) records potential changes on the scalp, originating from neurons in the gray matter [27]. These potential changes are typically captured by an electrode collection system positioned on the brain’s surface [21]. Analysis of EEG provides a comprehensive spectrum of physiological, psychological, and pathological insights [18]. However, owing to the high temporal resolution of EEG signals, they are susceptible to various complex noises, including cardiac artifacts, ocular artifacts, muscle artifacts, and external interference [10]. The persistent presence of these disruptive noises during data acquisition poses a substantial challenge in extracting pure EEG signals, imposing significant constraints on subsequent EEG research and applications [15]. Hence, an effective EEG denoising method is urgently required to mitigate the noise in acquired EEG signals while preserving essential information for further EEG research.
Numerous traditional denoising methods have been proposed for mitigating noise in EEG signals, including regression-based and adaptive filter-based approaches. Specifically, regression-based methods derive the noise signal using a predefined noise template. The computed noise signal is then employed as a suppression signal to eliminate noise from the EEG signal, yielding the denoised result [13] [5]. In contrast, the adaptive filter method embraces a distinct concept, suppressing noise by predicting filter coefficients directly from the input EEG signals [7] [11]. Nonetheless, conventional methods possess certain drawbacks. Notably, hyperparameter configuration in traditional methods significantly influences the efficacy of EEG noise removal, necessitating researchers’ empirical awareness to set reasonable parameters. Additionally, the traditional approach risks suppressing crucial EEG information while eliminating noise, potentially impacting subsequent research endeavors.
@@
The EEG signal is a complex waveform characterized by nonlinear features crucial for its analysis. Therefore, denoising methods must preserve these nonlinear features while eliminating noise [18]. The advancement in computer processing power and the expansion of EEG datasets [27] have spurred recent research endeavors to leverage deep learning for EEG signal denoising. Commonly employed architectures for EEG denoising networks include feedforward neural networks (FNN) [3] [25], convolutional neural networks (CNN) [1] [18], and recurrent neural networks (RNN) [26], along with their variations, such as long and short-term memory networks (LSTM) [14] [27]. As EEG is collected in the time dimension, establishing a temporal relationship between sampling points, basic network architectures have demonstrated significant improvement in denoising performance compared to traditional methods. However, they face challenges either in retaining temporal information or lacking global modeling capability while preserving temporal information. To address this, some studies have explored integrating the Transformer model [22] into EEG denoising tasks, as it [16] effectively preserves temporal information and enables efficient data parallel computation, yielding notable results.
@@
In recent years, with the rapid advancement of large-scale language modeling (LLM), a novel network called Retentive Network [19] has emerged. Retentive Network exhibits a favorable disposition towards temporal information, boasts robust global modeling capabilities for nonlinear features, and demonstrates commendable performance. However, when applied directly to denoise EEG signals, a challenge arises. This stems from the fact that EEG signals, akin to natural language, possess temporal characteristics and encompass global nonlinear features. Unfortunately, using Retentive Network directly for EEG denoising is unfeasible due to a misalignment in the dimensional requirements. Retentive Network, designed for two-dimensional input, conflicts with the one-dimensional nature of EEG signals. Unlike the approach in [16], reshaping a 1D signal into a 2D format results in a fixed sum of input dimensions after reshaping, compromising subsequent network feature extraction. To address this issue, we propose a signal embedding method capable of transforming the signal into a sequence of arbitrary length and embedding dimensions, enhancing network flexibility. Additionally, while EEGdenoiseNet [27] introduces a standard deep learning EEG dataset, expediting the development of EEG denoising methods, the dataset remains unprocessed, lacking sample pairs. This necessitates mixing various noise types (muscle artifacts and eye artifacts) during preprocessing. Divergent data preprocessing approaches may yield disparate network results, impeding the comparison of methodologies. To mitigate this, we curated an open-source dataset using the huggingface datasets library [12] from preprocessed data.  [^1] https://github.com/woldier/EEGDiR  [^2] https://huggingface.co/datasets/woldier/EEGDiRDataset
@@
    We introduce the Retentive Network, an advanced infrastructure in natural language processing, to the domain of EEG signal denoising. This integration not only offers a novel avenue for exploring this intersection but also broadens the scope of related research. Through this incorporation, we successfully establish an innovative approach for EEG signal denoising, thereby expanding the possibilities for further research in this domain.
@@
    To effectively integrate 1D EEG signals with our proposed network, we introduce a method called signal embedding. Unlike a simple reshaping of the signal into 2D, this method achieves an organic fusion of the 1D EEG signal and the network through a sophisticated embedding strategy. This innovative approach not only enhances the network’s adaptability but also contributes to an overall improvement in system performance.
@@
    When examining the standard deep learning EEG dataset provided by EEGDenoiseNet, we observe that the raw nature of the dataset and the absence of pairs of training samples hinder the comparison of different methods. To address this limitation, we create an open-source dataset using preprocessed data. This not only eliminates challenges related to noise and ensures data consistency but also facilitates the exploration of deep learning-based denoising methods for EEG signals.
@@ 2 Related work
The regression model serves as a pivotal tool for data smoothing and analysis in EEG signal processing [11]. The primary goal is to establish relationships, be they linear or nonlinear, between EEG signal channels and reference signals like EOG, EMG, or ECG channels. The overarching objective is to formulate mathematical equations through regression analysis to efficiently eliminate noise from EEG data [25]. Fundamentally, the method relies on a widely used regression-based approach, presupposing that each channel comprises a cumulative sum of pure EEG data and a fraction of artifact. This strategy is successful for channels with available reference data, utilizing exogenous reference channels such as ECG to eliminate various artifacts and improve the overall signal quality. However, it’s crucial to highlight a notable limitation: the regression-based method encounters challenges when handling electrooculogram (EOG) and electromyogram (EMG) signals due to the lack of an appropriate exogenous reference channel for these specific artifacts.
@@
The Wavelet Transform method [23] is employed to convert time domain signals into both time and frequency domains. This approach is preferred over Fourier transform due to its superior tunable time–frequency tradeoff and capabilities for analyzing non-stationary signals. The Wavelet Transform maps the signal to the wavelet domain, utilizing distinct properties and mechanisms of wavelet coefficients generated by signal and noise across different scales [4]. The primary objective of the method is to eliminate noise-generated wavelet coefficients while maximizing the retention of coefficients originating from real signals.
Traditional methods in EEG signal processing encompass regression modeling and wavelet transform methods. However, these methods exhibit drawbacks, particularly when processing complex EEG signals. The regression model may face limitations due to the absence of specific exogenous reference channels, and the wavelet transform method may lack sensitivity to the time-frequency characteristics of noise. To address these limitations, deep learning methods have emerged as an appealing alternative, leveraging their robust feature learning and representation capabilities, resulting in notable achievements in EEG denoising tasks.
@@
Figure 1: The diagram illustrates the training procedure of the deep learning method. The dataset comprises sample pairs, namely Noisy EEG signal and Noise-free EEG signal, representing EEG signals with and without noise, respectively. Throughout the training process, the Noisy EEG signal serves as the network input. The network, in turn, produces the Denoised EEG signal, which is utilized as the input for the subsequent training steps. The Loss Function calculates the disparity between the Denoised EEG signal and the Noise-free EEG signal, facilitating the optimization of the network. The black dashed box delineates the inference phase of the network, omitting the optimization component. This inference stage can be likened to the end-to-end output where the Noisy EEG signal is input into the network, yielding the Denoised EEG signal as the output.
In the training phase of the deep learning method, we utilize the framework illustrated in Figure 1, where Noisy EEG signal and Noise-free EEG signal form the pairs for training. The objective is to model the EEG signal with noise, aiming to produce an output that closely resembles the Noise-free state through the learning of network weight parameters. The training initiates with the input of the Noisy EEG signal into the network, generating the corresponding Denoised EEG signal. The disparity between this Denoised EEG signal and the actual Noise-free EEG signal is quantified as a loss, computed by the Loss Function. The black dashed box in the figure emphasizes the distinction between the signal and the Noise-free EEG signal.
The black dashed box in the figure delineates the network’s inference process. During inference, the network directly takes the noisy EEG signal as input and produces the denoised EEG signal as output, bypassing the optimization of weighting parameters. This end-to-end inference capability empowers the network to denoise new and unknown EEG signals in practical applications. The overarching objective of the entire training process is to achieve noise suppression by optimizing the network parameters, enabling it to extract genuine signal features amidst noise interference. Such a design facilitates the network in learning a more efficient representation, subsequently enhancing its denoising performance and providing robust support for real EEG signal processing.
Given the shared structure between the training and inference processes, researchers can concentrate their efforts on investigating the network’s architecture. This facilitates a more profound exploration of the application of deep learning in EEG signal processing. In recent years, deep learning methods have made significant strides in domains such as natural language processing [22] [20] and computer vision [6] [17]. Notably, their efficacy extends to signal processing, demonstrating impressive performance in signal denoising [18] [25] [27] [16].
@@ 3.1 Overall structure of the EEGDiR network
Figure 2: (a) illustrates the architecture of the EEGDiR network. This network generates hidden dimensions through Patch Embedding and obtains the output via linear projection and transformation following multi-level DiR Block processing. EEGDiR operates as an end-to-end model, taking a noisy signal as input and producing a noise-free signal, denoted as $x_{h}at$. The DiR Block, depicted in (b), comprises Pre-Norm, Multi-scale Retention, and Residual Connection, with Pre-Norm utilizing Layer Normalization. The SignalEmbedding structure, outlined in (c), involves segmenting the input sequence into new sequences based on the patch size. The hidden dimension after reshaping aligns with the patch size, and after linear projection, it matches the final hidden dimension.
[…]
  • Layer Normalization
    Layer Normalization (LN) [2] is another normalization technique that serves as an alternative to traditional Batch Normalization (BN) [8]. While BN can face challenges in performance when dealing with small batch sizes, Layer Normalization aims to address these issues by normalizing at the layer level. In contrast to BN, which computes mean and variance based on the entire batch, LN calculates normalization statistics independently for each layer. Specifically, the mean and variance are computed over all channels within a single layer, decoupling the normalization process from the batch size. This allows Layer Normalization to better adapt to varying batch sizes and perform effectively in scenarios where BN might struggle. By normalizing at the layer level, LN provides a robust estimation of statistics and can be advantageous in situations where small batch sizes or imbalanced data distribution pose challenges. This approach contributes to improved network performance and helps overcome limitations associated with traditional normalization methods under certain conditions.
  • Residual learning
    The use of residual learning is an effective technique to further enhance feature learning in deep convolutional neural networks (CNNs). Deeper CNN architectures have the capability to capture more abstract features and improve performance in tasks such as classification and regression. However, simply increasing the depth of a network does not always lead to better performance, as it can suffer from issues like vanishing or exploding gradients [6]. Residual learning, introduced in the context of the ResNet architecture, addresses these problems by incorporating residual blocks that allow for the direct propagation of information from earlier layers to later layers. By using skip connections, residual learning enables the network to learn residual mappings, which can help mitigate the degradation problem associated with increasing network depth. These residual connections provide shortcuts for gradient flow, making it easier for the network to optimize the deeper layers and improve overall performance. In addition to addressing gradient related issues, residual learning has been observed to enhance network performance. The skip connections allow for the preservation of important information from earlier layers, enabling the network to effectively learn complex representations. This approach has demonstrated improved convergence and achieved state-of-the-art results in various deep learning tasks.

### v2

@@ Abstract
Electroencephalogram (EEG) signals are pivotal in clinical medicine, brain research, and neurological disorder studies. However, their susceptibility to contamination from physiological and environmental noise challenges the precision of brain activity analysis. Advances in deep learning have yielded superior EEG signal denoising techniques that eclipse traditional approaches. In this research, we deploy the Retentive Network architecture—initially crafted for large language models (LLMs)—for EEG denoising, exploiting its robust feature extraction and comprehensive modeling prowess. Furthermore, its inherent temporal structure alignment makes the Retentive Network particularly well-suited for the time-series nature of EEG signals, offering an additional rationale for its adoption. To conform the Retentive Network to the unidimensional characteristic of EEG signals, we introduce a signal embedding tactic that reshapes these signals into a two-dimensional embedding space conducive to network processing. This avant-garde method not only carves a novel trajectory in EEG denoising but also enhances our comprehension of brain functionality and the accuracy in diagnosing neurological ailments. Moreover, in response to the labor-intensive creation of deep learning datasets, we furnish a standardized, preprocessed dataset poised to streamline deep learning advancements in this domain.
@@ 1 Introduction
Electroencephalography (EEG) measures neural activity as potential fluctuations on the scalp, primarily emanating from the brain’s gray matter [42] [26]. This neural activity is detected via an array of electrodes strategically placed on the scalp [35] [34]. Analyzing EEG data yields a broad range of physiological, psychological, and pathological insights [29]. Yet, the high temporal resolution characteristic of EEG signals renders them vulnerable to diverse and complex noise sources such as cardiac, ocular, and muscular artifacts, as well as environmental interference [14]. The prevalent intrusion of these noises during the acquisition phase significantly hampers the isolation of unadulterated EEG signals, thus severely restricting advancements in EEG-related research and practical applications [24] [33]. Consequently, there is an imperative need for a robust EEG denoising technique that effectively reduces noise without compromising critical signal information, which is vital for advancing EEG research.
A multitude of traditional denoising techniques for EEG signal enhancement has been advanced, encompassing both regression-based and adaptive filter-based methodologies. Regression-based strategies involve estimating the noise component with a pre-established noise model and subsequently subtracting this estimate to purify the EEG data, thereby producing a cleaner signal [21] [8]. Conversely, adaptive filtering operates on a fundamentally different principle, adjusting filter coefficients in real-time based on incoming EEG data to attenuate noise [10] [16]. However, these conventional techniques have their limitations. The tuning of hyperparameters in these methods critically affects the denoising performance, requiring expert judgment to optimize settings. Moreover, there is a risk of losing vital EEG information during noise reduction, which could detrimentally affect further analytical work.
@@
The EEG signal is a complex waveform characterized by nonlinear features crucial for its analysis. Therefore, denoising methods must preserve these nonlinear features while eliminating noise [29]. The advancement in computer processing power and the expansion of EEG datasets [42] have spurred recent research endeavors to leverage deep learning for EEG signal denoising. Commonly employed architectures for EEG denoising networks include feedforward neural networks (FNN) [4] [40], convolutional neural networks (CNN) [1] [29], and recurrent neural networks (RNN) [41], along with their variations, such as long and short-term memory networks (LSTM) [22] [42]. As EEG is collected in the time dimension, establishing a temporal relationship between sampling points, basic network architectures have demonstrated significant improvement in denoising performance compared to traditional methods. However, they face challenges either in retaining temporal information or lacking global modeling capability while preserving temporal information. To address this, some studies have explored integrating the Transformer model [36] into EEG denoising tasks, as it [25] effectively preserves temporal information and enables efficient data parallel computation, yielding notable results. In EEG signal processing, it is critical to accurately capture temporal order information, as patterns and features in these signals are often synchronized with specific neurophysiological events. Although the Transformer provides a powerful algorithm to process sequential data through an attentional mechanism, it does not explicitly encode the temporal order of the signals, but rather introduces this information indirectly through positional encoding. This may lead to insufficient sensitivity to the temporal dynamics of EEG signals in some cases.
@@
In recent years, with the rapid advancement of large language model (LLM), a novel network called Retentive Network [30] has emerged. Retentive Network gains its intuitive understanding of the temporal order in a sequence by introducing the decay mask Retention mechanism. With this mechanism, the input sequences are made naturally chronologically sequential, thus providing an effective way to model the inherent temporal dynamics of EEG signals. This approach is more suitable for processing EEG data as it is able to capture and utilize changes in bioelectrical activity over time, and thus may perform better than the standard Transformer model in practical applications. Retentive Network exhibits a favorable disposition towards temporal information, boasts robust global modeling capabilities for nonlinear features, and demonstrates commendable performance. However, when applied directly to denoise EEG signals, a challenge arises. This stems from the fact that EEG signals possess temporal characteristics and encompass global nonlinear features. Unfortunately, using Retentive Network directly for EEG denoising is unfeasible due to a misalignment in the dimensional requirements. Retentive Network, designed for two-dimensional input, conflicts with the one-dimensional nature of EEG signals. Unlike the approach in [25], reshaping a 1D signal into a 2D format results in a fixed sum of input dimensions after reshaping, compromising subsequent network feature extraction. To address this issue, we propose a signal embedding method capable of transforming the signal into a sequence of arbitrary length and embedding dimensions, enhancing network flexibility. Additionally, while EEGdenoiseNet [42] introduces a standard deep learning EEG dataset, expediting the development of EEG denoising methods, the dataset remains unprocessed, lacking sample pairs. This necessitates mixing various noise types (muscle artifacts and eye artifacts) during preprocessing. Divergent data preprocessing approaches may yield disparate network results, impeding the comparison of methodologies. To mitigate this, we curated an open-source dataset using the huggingface datasets library [17] from preprocessed data.  [^1] https://github.com/woldier/EEGDiR  [^2] https://huggingface.co/datasets/woldier/eeg_denoise_dataset
@@
    Proposal of Signal Embedding. In order to efficiently extract features from EEG signals, we introduce a method called signal embedding, which adds an embedding dimension to EEG signals. This method achieves the enhancement of feature information of EEG signals through the embedding strategy. The introduction of this method not only enhances the adaptability of the network, but also positively impacts the overall improvement of the system performance.
@@
    Introducing Retentive Network for EEG Signal Denoising. For the first time, we introduce the Retentive Network architecture into the field of EEG signal denoising to address the temporal nature of EEG signals, providing a new way to explore the intersection of EEG signals and natural language processing and expanding the scope of related research. It provides a new way to explore the intersection of EEG signal and natural language processing, and also expands the scope of related research.The introduction of Retentive Network allows us to take full advantage of its time-series information friendly and global modeling, thus realizing a new denoising method for EEG signal.
@@
    Provide open source datasets. When examining the standard deep learning EEG dataset provided by EEGDenoiseNet, we observe that the raw nature of the dataset and the absence of pairs of training samples hinder the comparison of different methods. To address this limitation, we create an open-source dataset using preprocessed data. This not only eliminates challenges related to noise and ensures data consistency but also facilitates the exploration of deep learning-based denoising methods for EEG signals.
@@ 2 Related work
2.1 Traditional Methods
Regression methods typically depend on exogenous reference channels such as EOG, EMG, or ECG [40] to model and eliminate associated noise [16]. The efficacy of these methods is contingent upon the quality of the reference channels. Poor quality or absence of these channels significantly undermines the performance of the regression model. Furthermore, many regression methods, especially linear regression, presuppose a linear relationship between the data. However, the association between EEG signals and noise is often nonlinear, particularly when noise sources are complex, such as muscle activity or eye movement. This complexity necessitates more sophisticated nonlinear models for effective noise removal.
@@ 2.1 Traditional Methods
The Wavelet Transform method [38] is used to convert time-domain signals into time and frequency domains. This method is favored over the Fourier Transform due to its better tunable time-frequency tradeoff and its capability to analyze non-stationary signals. It operates by mapping the signal into the wavelet domain, where distinct properties of wavelet coefficients generated by signal and noise at various scales are utilized [5]. The primary goal is to eliminate noise-generated wavelet coefficients while preserving those from the actual signals. However, this method may lack sensitivity to the specific time-frequency characteristics of the noise in complex EEG signals.
To overcome these challenges, deep learning methods have become a promising alternative, thanks to their robust feature learning and representation capabilities, achieving significant successes in EEG denoising tasks.
@@
Figure 1: The diagram illustrates the training procedure of the deep learning method. It includes the Data Processing stage, where raw EEG data and various artifacts undergo preprocessing to create a suitable dataset for deep learning. The dataset comprises sample pairs, namely Noisy EEG signal and Noise-free EEG signal, representing EEG signals with and without noise, respectively. Throughout the training process, the Noisy EEG signal serves as the network input. The network, in turn, produces the Denoised EEG signal, which is utilized as the input for the subsequent training steps. The Loss Function calculates the disparity between the Denoised EEG signal and the Noise-free EEG signal, facilitating the optimization of the network. The black dashed box delineates the inference phase of the network, omitting the optimization component. This inference stage can be likened to the end-to-end output where the Noisy EEG signal is input into the network, yielding the Denoised EEG signal as the output.
2.2 Deep Learning Methods
In recent years, deep learning has significantly advanced in fields such as natural language processing [36] [32] [23] and computer vision [9] [27]. Notably, its application to signal processing has demonstrated remarkable efficacy in signal denoising [29] [40] [42] [25] [37] [40] [2].
@@ 2.2 Deep Learning Methods
2.3 Dataset Preparation
In order to prepare data for deep learning, the raw EEG data and various artifacts undergo preprocessing, as illustrated in the left half of Fig. 1. This includes cropping, signal stacking, normalization, and other operations to create a dataset suitable for deep learning. The creation of this dataset is foundational to our study, as it provides immediate access to preprocessed EEG data for training deep learning models. However, data preprocessing is laborious and time-consuming, underscoring the importance of providing an out-of-the-box (i.e., no data preprocessing required) dataset readily available to researchers.
During the training phase of the deep learning method, employ the framework depicted in the right half of Fig. 1, where noisy EEG signals and noiseless EEG signals form pairs for training. The objective of training is to model the noisy EEG signals to produce outputs closely resembling the noise-free state by learning the network’s weighting parameters. Initially, the noisy EEG signal is inputted into the network to generate the corresponding denoised EEG signal. The discrepancy between this denoised EEG signal and the actual noise-free EEG signal is quantified as a loss, calculated by the loss function.
The black dashed box in the figure delineates the network’s inference process. During inference, the network directly takes the noisy EEG signal as input and outputs the denoised EEG signal, bypassing the optimization of the weighting parameters. This end-to-end inference capability enables the network to denoise new and unknown EEG signals in practical applications. The primary objective of the entire training process is noise suppression by optimizing the network parameters to extract true signal features amidst noise interference. This design facilitates the network to learn more effective representations, enhancing its denoising performance and providing robust support for real EEG signal processing.
Given the shared structure between the training and inference processes, and provide an out-of-the-box dataset readily available to researchers. Researchers can concentrate their efforts on investigating the network’s architecture. This facilitates a more profound exploration of the application of deep learning in EEG signal processing.
@@ 3.1 Overall structure of the EEGDiR network
Figure 2: (a) illustrates the architecture of the EEGDiR network. This network generates hidden dimensions through Signal Embedding and obtains the output via linear projection and transformation following multi-level DiR Block processing. EEGDiR operates as an end-to-end model, taking a noisy signal as input and producing a noise-free signal, denoted as $\hat{x}$. The DiR Block, depicted in (b), comprises Pre-Norm, Multi-scale Retention, and Residual Connection, with Pre-Norm utilizing Layer Normalization. The Signal Embedding structure, outlined in (c), involves segmenting the input sequence into new sequences based on the patch size. The hidden dimension after reshaping aligns with the patch size, and after linear projection, it matches the final hidden dimension.
In this paper, we present EEGDiR, a novel network model tailored for EEG signal denoising. Our model incorporates Retentive Network into the realm of EEG signal denoising, introducing innovative perspectives to signal processing tasks. The overall network structure (see Fig. 2(a)) involves processing the noisy signal y through signal embedding, elaborated later, to augment the embedding dimension. Subsequently, the noise signal y undergoes processing via stacked DiR Blocks at multiple levels, with the final output linearly projected to match the input dimension. EEGDiR operates as an end-to-end model, taking a noisy input signal y and generating the corresponding noiseless signal $\hat{x}$. Fig. 2(b) depicts the structure of the DiR Block, which begins with pre-Norm, followed by multi-scale Retention and a Residual Connection [9]. By using skip connections, residual learning enables the network to learn residual mappings, which can help mitigate the degradation problem associated with increasing network depth. These residual connections provide shortcuts for gradient flow, making it easier for the network to optimize the deeper layers and improve overall performance. The output of the residual join undergoes pre-Norm once more before serving as the input for the Fully Functional Network (FFN). The term "pre-Norm" refers to Layer Normalization, employed for normalization before each submodule. Layer Normalization (LN) [3] is another normalization technique that serves as an alternative to traditional Batch Normalization (BN) [12]. While BN can face challenges in performance when dealing with small batch sizes, Layer Normalization aims to address these issues by normalizing at the layer level. It is crucial to note that an FFN typically includes a fully-connected module with a hidden layer that doubles the hidden dimension. The FFNs output layer reduces the hidden dimension to align with the input. The dimensions of input and output vectors remain constant across DiR and its submodules. Figure 2(c) illustrates the signal embedding structure, the input sequence is segmented into new sequences of length $l_{s}//patch\_size$ based on the patch size. The initial hidden dimension of these sequences is equal to the patch size after reshaping. Through linear projection, the hidden dimension of the projected sequences matches the final hidden dimension. Given that EEG signals may encompass multiple frequencies and require effective feature capture at different scales, signal embedding is introduced to intelligently handle temporal information at varying scales. It enhances context preservation and temporal relationship retention by merging consecutive samples into a patch.